UT Apr 17 2022 - DEC Night




Night Info

  • Moonrise 18:59
  • Sunset 18:50
  • Nautical Twilight Ends 19:46
  • Astro Twilight Ends 20:17
  • Astro Twilight Starts 04:21
  • Nautical Twilight Starts 04:52
  • Sunrise 05:48
  • Moon illumination 0.996

Staff

  • OSA: J.Williams
  • ISp: C.Bolyard/Mtn
  • Mtn.Manager: J.Prothro
  • ISA/Observer: B.Rothberg, R.T.Gatto
  • Special Guests:


Night Plan

  • Starting config bino-LUCIs, authorize LUCI-1 for AO
  • LUCI-1 AO/ESM
    • DONE OSCO w/NCPA on - 0.5hr - 19:40-20:10
    • DONE ESM Bin 2 testing - 1hr
      • seeing up to 1.5", thin clouds (just make sure it closes at Bin 2)
      • No need for SL data in parallel - trying to rail camera so scripts may not execute fully
      • PA 0,60,120,180,240,300 to test railings
        • M67 AO Ref R=12.2 but set to 12.6 to force Bin 2 - visible until 23:45hr
        • M92 AO Ref R=11.3 but set to 12.6 to force Bin 1 - visible 23hr-twilight
    • DONE ESM Bin 4 test
    • Close Separtion Targets test (0.1"-2") - 1-2hr
      • in OT: AO STARS - FULL DITHERS *NEW*-->Binaries
        • Target 2 BD+00 2888 R=10.7 - separation 0.225" - need 256x2048 subframe - visible twilight-02:30hr
          • Target 5 GD 135 R=14.6 - separation 0.3" - no subframe - visible twilight-01:30hr
        • Target 4 SBSS 1148+564 R=14.9 separation 0.6" - no subframe - visible twilight - 04hr
        • Target 6 2MASSJ1128+3136 R=13.5 - 1" separation - no subframe - visible twilight - 03hr
        • Target 7 2MASSJ13061537+2043444 R=8.5 - 1.6" separation -64x2048 subframe - visible twilight-twilight
          • Note: given separation and lack of GS only 1 PA script is available - companion should be visible directly below main star
          • Target 9 2MASSJ10590395+6349283 Bin 2 R=12.7 - 1.57" separation - No subframe backup - visible twilight-03:30hr
        • Target 3 BD+33 2137 R=10.3 - 2.4" separation no subframe
          • Target 10 2MASSJ15553957+4025135 Bin2/3 R=13.3 - 2.3" separation - visible 21:45-twlight
    • Faint Limit Test WFS (Rmag>16.5)
      • AO Stars FULL DITHERS NEW folder
        • RA08-09hr Star 10 R=16.6 (17.8) - twlight-00hr
        • RA10-11hr Star 5 R=16.6 (17.8) - twlight-01:45hr
        • RA10-11hr Star 6 R=17.0 (18.2) - twilight-03hr
        • RA12-13hr Star 4 R=16.6 (17.8) - all night
        • RA12-13hr Star 8 R=17.3 (18.5) - 21:15-04hr
        • RA12-13hr Star 9 R=16.7 (17.9) - 20:30-04:30hr
        • RA14hr-15hr Star 4 R=17.1 (18.3) -20:30-twilight
    • Interesting Targets (if time)
      • Cat's Eye Nebula - 23hr-twilight
        • FeII and Hbeta needed (H2 already collected)
        • Dec +66
      • Planetary Nebula BD +303639
        • H2,FeII,Hbeta, Ks - 23hr twilight
        • Dec +30
      • DDO82 - dwarf galaxy w/Bin 1 off-axis star
        • J,K - twilight-03:30hr
        • Dec +70
  • Reconfigure for LBCs by 00:30
    • LBC Extra Focal images in equilibrium -IT 8608, IT 8592 - 1.5hr
    • PHAs recovery - 1hr
      • CP for collimaion is at 14:39 +34:28 - target is at 14:05 +33:52 tonight - this is close enough
  • Reconfigure for MODS-Bino for D-time science by 03:45hr
    • V1674 Her Followup - MODS Bino - M.Wagner - 0.75hr
    • Spectrophotometric star (if clear take all modes) - BD+33

Poor Weather Backup

Eng Tasks
  • Poor weather ESM - M67/M92 - track performance in poor weather
  • Mixed-mods on-sky LUCI+MODS test -0.5hr
    • send active bino preset
  • DONE LUCI Photometric Standards (any seeing, if so bad can't guide, use track mode) - requires clear skies
    • /home/lbto/Service_Observing/2022A_D_E/PoorWeather/LUCI_Phot_Standards
    • Select stars based on visibility, need to get stars at two different airmasses
  • Pointing Model - DX needs improvement after ASM re-install
    • Seeing up to ~2.5"
    • Use LUCIs
      • N375 cameras, N30_FieldStop (aka the "Pointing") masks, K-band, 2.51 sec NDIT=2
      • no flexure compensation (doesn't matter with the pointing mask in)
      • Before taking pointing data, take at least 2 images with dROT=180 for rotator center (I usually do 4)
      • use PtStars in IRAF - current or near current LST - re-run every 2-4 stars with current LST
      • active preset to collimate on-axis, park probe, then take image - 1st image
      • turn PCS logging ON (it is possible this is always on now...check)
      • send track mode presets to stars after that Note: if you use off-axis stars and let it collimate at each point, you can also make a collimation model
      • Turn PCS logging off, note filename
Science
  • Rothberg - Local Mergers
    • OT: /home/lbto/Service_Observing/2022A_D_E/PoorWeather/Rothberg/LocalMergers/COBandhead_Mergers_V43.xml
    • Not in order of priority
      • NGC 4004 - 21hr-twilight Dec +27
      • NGC 3597 - 23hr-02hr Dec -23
      • DONE NGC 6598 - 03hr-twilight Dec +69
      • UGC 9829 - 01:30hr-twilight Dec -01
      • DONE UGC 10607 - 02hr -twilight Dec +26
      • UGC 10675 - 02hr-twilight Dec +31


Night Log

Closed Dome Prep

  • UT 00:02 - starting up LUCI GUIs
    • init all, test darks L1 0001-0002, L2 0001-0002
    • check N30fieldstop on LUCI-1
    • camera filter wheel error moving from N375 to N30
    • init at N375
    • moved to N180 OK
    • then moved to N30 not ok
    • says -354 at N375, moving -500 motor steps to hit neg limit for N1.8, can't seem to turn on the neg limit swithch - stopped at -1234 - but HIRAMO says we are there (is neg limit switch not working?)
    • moving to N30, seems to have gotten to there, HIRAMO shows switch 11 and 12 on
    • Fieldstop check L1 0003-0006
  • UT 00:15 - starting LBC-Red
  • UT 00:33 - taking LBC Red biases - some banding first bias, 2nd is ok
  • UT 00:41 - O.Kuhn checking zernike insertion script for primary mirrors, B.Rothberg waking MODS on OBS2
  • UT 00:46 - taking simSnap - all OK

On-Sky

  • UT 01:56 - opening enclosure - will do an OSCO first
  • UT 01:59 - left shutter doesn't want to open
  • UT 02:10 - troubleshooting left shutter
  • UT 02:45 - still troubleshooting left shutter
  • UT 03:12 - still troubleshooting left shutter
  • UT 03:33 - attempting to open both shutters
  • UT 03:36 - left alarm 1002 failed run alarm - clearing it in ECS works briefly, then it appears again
  • UT 03:41 - looks like both sides ok now - doors opening
OSCO
  • UT 03:45 - sending pointing pointing preset
  • UT 03:54 - collimation preset shows 1.3" in guiders both sides
  • UT 03:59 - J.Williams sending OSCO script - BD+70 622
    • Rmag=7.48 - 1.6"-.7"on SX Guider
    • CCD47 offline
  • UT 04:03 - resending preset
    • failed again - no connection w/message daemon
    • Sun Apr 17 04:04:14.996 2022 left Preset alert for LUCIFER mode ADAPTIVEACE_ACTIVE: PresetAO {FLAO}[849] - [AOException] presetAO: FATAL - thGetVar ttctrl.L.ENABLEOFFSET.REQ error(-3606): No connection with MsgDSun Apr 17 04:04:15.018 2022 left Preset command failed for LUCIFER mode ADAPTIVEACE_ACTIVE with guidestars, offset, AO reference, Sound=presetFailed.wav
  • UT 04:05 - Guider shows 1.43" seeing, DIMM 1.45"
  • UT 04:09 - JCGuerra working on issue - seeing is osciallatin 1.3"-2.1" - reseting BCU, CCD47, bayside stages
  • UT 04:10 - CCD47 is live again - but bayside stage won't initialize at 184 instrument rotator angle - J.Williams is moving to 341 deg - JCGuerra will reset bayside stage and then we'll resend the preset
  • UT 04:15 - resending preset
  • UT 04:17 - CCD47 is not live again, BCU timeout, guider seeing is 1.1" DIMM 1.3"
    • Sun Apr 17 04:18:52.511 2022 left Preset alert for LUCIFER mode ADAPTIVEACE_ACTIVE: PresetAO {FLAO}[852] - [AOException] presetAO: FATAL - Reading shared buffer mirrorctrl.L:TECHVIEWER_65536 (-5001) TIMEOUT_ERROR
      Sun Apr 17 04:18:52.531 2022 left Preset command failed for LUCIFER mode ADAPTIVEACE_ACTIVE with guidestars, offset, AO reference, Sound=presetFailed.wav
  • Might need some time to fix - seeing

LUCI Photometric Standards

GD153
  • GD153 - bino phot standard - last seeing was 1.5"
  • UT 04:24 - sending preset
  • UT 04:27 - missed on right, need to add back IE CA for bino
  • UT 04:31 - seeing is 1.2"-1.5" - JCGuerra suggested we can try again just as we sent script - ran some diagnostics
    • Airmass 1.15-1.12
    • skipped z-band
    • L1 0007-0025
    • L2 0003-0018
  • UT 04:45 - done

OSCO Take II

  • UT 04:45 - authorizing LUCI-1 only
  • UT 04:49- light in CCD47 now - guider show 1.2"-1.4"
  • WFSMag = 9.1, Rmag = 7.48
  • Seeing blew up to 2.1" on DIMM - bit on mirror in OCAM viewer
  • UT 04:55 - seeing back down to 1.2" on DIMM
  • JC Guerra - checking to see what high order gains triggers the bite
  • UT 04:56 - seeing bliew up again - now 1.7"
  • L1 0023 - no NCPA
  • L1 0024-0042 w/NCPA
    • Tracking Numbers first position
    • => Command line arguements processed
      Set 0 Tracknum 20220417_050002 LuciName unknown
      Set 1 Tracknum 20220417_050015 LuciName unknown
      Set 2 Tracknum 20220417_050026 LuciName unknown
      -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220417_050002.log
  • UT 05:03 getting skipframes
    • Tracking Numbers 2nd position
    • w/NCPA

      [telescope@obs1 ~]$ ./get_ao_telemetry

      => Command line arguements processed
      Set 0 Tracknum 20220417_050248 LuciName unknown
      Set 1 Tracknum 20220417_050300 LuciName unknown
      Set 2 Tracknum 20220417_050311 LuciName unknown
      -> wrote file /home/AOeng/doug/get_telemetry/OSCO/20220417_050248.log
  • UT 05:04 - finished w/1.7" on DIMM

BIN 2 ESM Test

  • L1 0043
  • UT 05:05 - M67 - Bin 2 R=12.2 (R=12.6 forced) - PA = 0, Ins Rotator = 313.8
  • WFSMag=14.06
  • UT 05:09 - did activate TSS
  • OBS5 brieflt froze
  • GCS guider 1.3"
  • UT 05:13 - and then seeing blewup to 1.8"-2"
  • UT 05:17 - taking a while, was running GOpt - but this is ESM
  • getting skipframes
  • UT 05:19 - M67 PA=60, Ins Rotator = 373.2
  • Bayside stage didn't move - star is found on CCD47 - got stuck for a while
  • UT 05:24 - getting torques on elevation drive, may need to move off
  • UT 05:25 - aborted image

Bin 4 Test - M51 - L1 ESM only

  • UT 05:29 - sending preset PA=90, Ins Rotator=216.3
  • DIMM 1.7"
  • WFSMAG=15.72, AORef=15.4 (was sent as 16.51)
  • JC.Guerra checking on right gains to use - will run through some of script
  • UT 05:37 - DIMM 1.4"
  • UT 05:39 - not taking data - no offset was sent by LUCI thinks it was, skip to line and resent offset
  • first image 0.57" FWHM, 3rd exposure 0.61"
  • UT 05:46 - DIMM 2"
    • L1 0044-0052
  • UT 05:54 - authorizing Bino

Bin 1 Test ESM/SL Mixed Mode M61

  • UT 05:56 - sending preset
  • UT 05:59 - seeing 1.6" on SX guider - DX started out way out of collimation (did this earlier as well)
  • star is in CCD47
  • centerstar is done, but nothing in OCAM for center pupils, not getting frames
  • taking another OCAM dark, PA=150, instrument roatator = 154.7
  • UT 06:10 - power bump? Azimuth drive conked out - lost preset - torques on AZ drive were high
  • UT 06:14 - sending new version with AORefstar = 14.52 - to see if this explains why can't see it on OCAM, Instrument Rotator=151.8
  • UT 06:16 - er....reset preset but IIFGUI is hanging on waiting for AZ and Rot - ut we aren't moving
  • UT 06:20 - resending preset - hydrostatic pressures increased on AZ drive
  • UT 06:21 - didn't find guidestars...doing a pointing check
  • UT 06:26 - re-sending ye again the preset
  • UT 06:28 - scrambled image on DX WFS - first two frames appeared scrambled - then OK
  • Checking to see if AORefstar is on OCAM
  • Guiders show 2" seeing
  • Rmag=10.67, WFSMag=15.7 (seems more likely UCAC5 was wrong)
  • first exposure 0.7", 2nd 1.07"
  • fell back to active preset - #055 - but according to OT - the AO Ref should have been near the center of the patrol
  • resending back to 3rd position - check to see if AO Ref star is on OCAM
  • M61_AO_PatrolFIeld_RefPosition_3rddither.png
  • AO bayside stages moved
  • #057 - 0,0 absolute offset
  • UT 06:47 - preset fell back to guide - loop opened - could be the seeing
  • UT 06:48 - resending the preset - 2.6" on SX guider, 2.4" on DX guider
  • UT 06:52 - JC Guerra trying to close loop on 2.5" seeing
  • loop closed but it fell back to a guide preset pretty quickly (likely because there's fizz in the AGW!) - oscillating between closed and open loop
    • L1 0053-0058
    • L2 0019-0024
  • UT 06:57 - we are more than done here - cancelling preset

Photometric Standards Part Deux

GD153
  • UT 06:58 - returning to GD153
  • Airmass 1.02
  • guiders show 2.47",2.2"
    • L1 0059-0078
    • L2 0025-0044

Rothberg Local Mergers Backup Science

UGC 10607
  • UT 07:18 - sending preset to telluric for UGC 10607
  • UT 07:25 - pixmap RTD1 had to reset stop start
    • L1 -10.3817",-2.045"
    • L2 -8.248", -1.4125"
    • DIMM 2.2", guiders show 1.4"-1.5"
  • UT 07:29 - staring spectra
    • L1 0079-0084
    • L2 0045-0050
  • UT 07:34 - sending preset to UGC 10607
  • missed on SX
  • UT 07:37 - pointing check
  • UT 07:40 - resending preset
  • 1.92", and 1.7" on SX and DX guiders
    • offsets
    • L1 -0.2144", -1.9625
    • L2 1.334", -1.3792"
  • UT 07:47-07:49 - had to restart RTDs - resetl stop, start - due to wacky color palette
  • same issue as last night with RTD2 changng into RA Dec not X and Y
  • >>> Command.exec() : execute "sync "...
    *** => NAXIS1=2048 NAXIS2=2048 BITPIX=32 => size=16777216
    *** => BZERO = 0.0
    *** => BZERO = 0.0 BSCALE = 1.0
    *** => Reading and autocutting in 0.053 s => 301.89 Mbyte/s
    *** Original pixels saved in cache [/home/lbto/.aladin/.aladin.32642/f63.tmp]
    *** Original pixels RAM free for luci2.20220417.0054[0]
    *** Plane [luci2.20220417.0054] load&build in 239ms
    Display FITS file luci2.20220417.0054 executing
    cview luci2.20220417.0054; cm auto BB noreverse
    [cview luci2.20220417.0054]...
    >>>> Command.exec() : execute "cview luci2.20220417.0054"...
    [cm auto BB noreverse]...
    >>>> Command.exec() : execute "cm auto BB noreverse"...
    Display FITS file luci2.20220417.0054 done
    Wait for Aladin plane luci2.20220417.0054 done
    Remove MOS planes executing
    rm CTR_refslits; rm CTR_reftargets; rm LMS_refslits; rm LMS_reftargets
    [rm CTR_refslits]...
    >>>> Command.exec() : execute "rm CTR_refslits"...
    [rm CTR_reftargets]...
    >>>> Command.exec() : execute "rm CTR_reftargets"...
    [rm LMS_refslits]...
    >>>> Command.exec() : execute "rm LMS_refslits"...
    [rm LMS_reftargets]...
    >>>> Command.exec() : execute "rm LMS_reftargets"...
    Remove MOS planes done
    FITS header keyword MRSLITNO not found in luci2.20220417.0054
    SelectedTwo: luci2.20220417.0052
    ReputTwo: luci2.20220417.0052
    SelectedTwo: luci2.20220417.0051
    ReputTwo: luci2.20220417.0051
    *** Original pixel sub-image (768,512,512,1024) extractedfrom cache [/home/lbto/.aladin/.aladin.32642/f63.tmp]
    >>>> PlanImage.recut(-209172.0,473205.0,false) on [768,512-512x1024] => min=-209172.0 max=473205.0
    *** Original pixel sub-image (922,937,237,237) extractedfrom cache [/home/lbto/.aladin/.aladin.32642/f63.tmp]
    >>>> PlanImage.recut(-263.91460000000006,64791.028999999995,true) on [922,937-237x237] => min=-263.91460000000006 max=64791.028999999995
  • UT 07:53 - starting spectra
  • Seeng on guiders 2", 2.6" - but the blobs in the middle of the galaxy are int he slit
  • we'll likely sit on this for a while given the seeing
  • setting a pause after 2nd dtither - will check on seeing in guider
  • hit continue when it came down below 2"
  • UT 08:54 - seeing is still highly variable - but the mins have dropped down to as low as 1.3"
  • UT 09:03 - El drive failed - occured in alast 90 sec of final dither - aborted exposure
    • might be a PLC problem
    • L1 0085-0099
    • L2 0051-0065
    • got 11 x 300sec x 2
  • UT 09:12 - putting LUCIs in safe config (mirror, blinds, no flexure) while J.Williams works on problem
    • unstable power?
  • UT 09:14 - just lost azimuth drive - restarred ok - elevation drive still stuck
  • UT 09:21 - OBS5 X2GO crashed - or someone stole it, J.Williams calling D.Rapozza for El drive issue - J.Williams heading up to telescope
  • UT 09:41 - power-cycleing TPLC seems to have fixed the El drive

Photometric Standards Part Trois

  • UT 09:44 - will do a pointing and collimation check nearby
  • UT 09:52 - collimation check shows 1.4"-1.9" seeing ... or maybe 2"
  • UT 09:54 -sending GD153 preset
  • Airmass 1.42-1.51
  • Seeing continues to be highly variable - and continues to be variable
  • UT 10:06 - J-band images 1.1"-1.4" FWHM
  • UT 10:09 - FWHMs on guiders jumping to 2.4"-2.7"
  • UT 10:12 - LUCI-1 z-band FWHM 2.39"
    • L1 0100-0119
    • L2 0066-0085

Rothberg Local Mergers Backup Science

NGC 6598
  • Continue yesterday's poor weather target
  • UT 10:14 - sending preset telluric for NGC 6598
  • UT 10:19- now the GS for the telluric show 1" FWHM?
    • L1 -8.88", 1.9036"
    • L2 -8.3479", 3.00"
    • DIMM shows 1.6", guiders show 1.1"
  • UT 10:24 - RTD1 showing RA Dec now too
  • UT 10:25 - starting spectroscopy
  • UT 10:29 - powering off LBCs - too much overhead to try and switch to them now
    • L1 0120-0125
    • L2 0086-0091
  • UT 10:31 - sending preset to science target NGC 6598 - let's see what seeing is now
    • L1 -0.566", -2.7377"
    • L2 -0.4236", -2.2503"
  • UT 10:44 - starting spectra
  • LUCI-2 countdown was 4000 seconds, manually aborted and started exposure manually
  • 2022.04.17 10:44:54 s | error high | luci.luc | Luci TWO | TWO_GEIRSServer | RMIGEIRSServiceImpl.java#checkOkAnswer(945) | GEIRS received response with error: ERROR 19 Command 'crep 1' returned errorcode = 19: (E_invcamera=19) invalid ROE response
    .
    2022.04.17 10:44:54 s | error medium | luci.luc | Luci TWO | TWO_ReadoutMana | GEIRSClient.java#setNDIT(121) | Could not commit value "1" by "setNDIT" to GEIRS from "rmi://localhost:60002/TWO_GEIRSServer"!
  • Seeing again is all over the place - one seconds its 1.53", the next its 1.2"
  • UT 11:03 J. Guerra powering off WFS and warming OCAM
  • UT 11:18 - now seeing seems consistently better - but too late to reconfigure for MODS as astro twilight is in 15min
  • UT 11:22 - putting MODS to sleep
    • L1 0126-0141
    • L2 0092-0107
  • UT 11:57 - all done, putting in safe config in prep for imaging cals
  • Tonight's DIMM
  • DIMM.png

Closed Dome Calibrations

  • N375 zJHKKs flats
  • L1 0142-0191
  • L2 0108-0117
  • UT 12:44 - LUCI-2 safe config
  • UT 12:45 - LUCI-1 safe config

-- BarryRothberg - 15 Apr 2022
Topic revision: r9 - 17 Apr 2022, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback